OpenStudy (anonymous):

30 men working 7 hrs a day can do a piece of work in 18 days. In how many days will 21 men working 8 hrs a day to do same work ?

OpenStudy (anonymous):

22.5 days equivalent to 23 days

OpenStudy (anonymous):

30 7 18 21 8 ? now 30 -> 7= 18 1 -> 7= 18*30 1 -> 1= 18*30*7 21 -> 8= 18*30*7/21*8 ==>45/2=22.5

OpenStudy (anonymous):

gud ans yaar! but in detailed ?

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

a/b+c/d=b/a+d/c,then prove that a3/b3+c3/d3=b3/a3+d3+c3.r u konw this answer?

OpenStudy (anonymous):

no buddy, i dnt know it !!

OpenStudy (anonymous):

30 men works 7 hours daily implies that 30*7 = 210 hrs daily. In 18 days, that would be 18*210 = 3780 hrs to complete the work. Now, 21 men works 8 hrs daily implies that 21*8 = 168 hrs daily. To complete the work now, will take 3780/168 = 22.5 days.

OpenStudy (anonymous):

great !!

OpenStudy (anonymous):

u want what u want to write with this solotion.sorry

OpenStudy (anonymous):

i dnt get u ?? thesujoysaha !!